# Slice-aware Resource Allocation and Admission Control for Smart Factory Wireless Networks

### v2

@@ III-D Admission Control Strategy - Rate Readjustments
The objective of our optimization problem is to minimize power usage while maintaining the target rate requirements for all users in each network slice. When transmission power requirements exceed the BS available power due to high target rates, it is necessary to readjust the rates of users in the CL slice to ensure high-priority tasks in the URLLC and TS slices can continue without interruption.
Each user is associated with a Lagrange multiplier, $\lambda_{i}$ during optimization as in algorithm 1. $\boldsymbol{\lambda}$ indicates how much the overall power consumption will change if users rate constraints are relaxed or tightened. We employ the rate readjustment algorithm detailed in [15], to proportionally lower the target rates across users using their multipliers. The algorithm adopts the well known Newton-Raphson method [16] to iteratively calculate the difference between available and required power based on $\boldsymbol{\lambda}$, efficiently reducing rates step by step. The adjustments continue until power consumption is just below the maximum allowable limit, managed by a tolerance value.
The proportional rate reduction $\Delta r(i)$ for each CL slice user $i$ is given by:
$\Delta r(i)=(p_{\text{opt}}-1)\left(\frac{\lambda_{i}}{\sum_{i=1}^{N}\lambda_{i}}\right),\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{C}$ (18)
where $p_{\text{opt}}=p_{\text{required}}/p_{\text{available}}$. The rate update formula is thus given by:
$r_{i}=\max\left((r_{i}-\Delta r_{i}),0\right),\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{C}$ (19)
By updating the rates with (19), users in CL slice contributing more to the power excess (higher $\lambda_{i}$) will experience a larger rate reduction. Thereby efficiently reducing the total power by targeting the reductions more strategically based on users contributions to the power overshoot.
@@ IV-A Scenario Description
For our simulation, we consider an indoor smart factory environment, characterised by densely packed reflecting elements of diverse shapes and sizes, and different user terminals each with varying QoS requirements that generate different traffic patterns. We consider a 5G BS deployed in this environment with a coverage area radius of 100 m having three slices, one each of CL, URLLC and TS slices. We use the 3GPP [17] non-line-of-sight (NLOS) path loss model for indoor factory (InF) dense clutter, low BS (InF-DL) scenario given in (20) in combination with Rayleigh fading, to obtain the channel gain in the links between the users and the BS:
@@ IV-B Slice Isolation and QoS Assessments
To evaluate the achievement of effective slice isolation, we show the mean user rates across the three slices in Fig.2. From this figure, we can observe that the mean rate of users in the CL slice experiences abrupt increases and decreases in response to the sudden changes in the number of users, decreasing and increasing respectively, at those specific times. However, the mean rate for the users in the URLLC and TS slices remains consistent, indicating complete isolation between the slices.
Figure 2: Mean rate of users per slice
Given that all QoS constraints where translated into rate constraints, meeting the target rates for each slice equates to fulfilling their respective QoS requirements. Fig.3 shows the sum rates of users in each slice. This figure clearly indicates that all the target rates per slice, as outlined in Table II, have been met, thereby demonstrating the achievement of the delay, jitter, reliability and bitrate constraints for all users. An initial transient for the sum rate in the CL slice is observed as a result of the transient behaviour in the control loop in (17).
Figure 3: Sum rate of users per slice
@@ IV-C Admission Control
Fig.4 shows the total BS power without admission control. We can observe that the power required to satisfy the QoS for all users across the three slices exceeds the available 23 dBm even when the number of users in the CL slice decreased at $6633>t<$ 66. This insufficiency indicates that the requirements of all active users will not be met, which can disrupt critical transmissions. To prevent this, we implement admission control to limit the required power to 23 dBm by selectively reducing the rates of users in the CL slice with less stringent requirements. The effects of this strategy are depicted in Fig.5 and Fig.6.
The total BS power when admission control is active is shown in Fig.5. The power requirement remains within the available limit of 23 dBm, confirming the success of our strategy. The quick drop in the total power observed at $t=33$ corresponds to the sudden decrease in the number of users at that time slot. Additionally, Fig.6(a) illustrates the rate adjustments in the CL slice during admission control. Notably, at $t<33$ when the required power initially exceeds 23 dBm, the sum rate of the CL slice is reduced from the target of 270 Mbps to approximately 259 Mbps. At $6633>t<$ 66, as the number of users in the CL slice decreases from 10 to 7, the sum rate is readjusted to around 267 Mbps. Furthermore, when the number of users increased from 7 to 12 at $t>66$, the sum rate reduced again to 251 Mbps.
@@
Figure 4: BS transmit power without admission control
Figure 5: BS transmit power with admission control
Figure 6: Sum rates (a) and mean rates (b) per slice during admission control
@@ V Conclusion and Future work
This research lays the groundwork for developing a dynamic, efficient, and autonomous slice-aware radio resource management system for smart factories. Future work will delve into the slicing resource allocation challenges posed by mobile terminals like AGVs and evaluate performance using realistic industrial scenarios with actual factory channel traces and its application to uplink communications.
@@ VI Appendix I
Take a sequence $\mu_{1}\leq\mu_{2}\leq\dots\leq\mu_{N}$ of real values, and define $F(\mu)$ a function to be minimized with respect to $\mu$:

### v3

@@ III-D Admission Control Strategy - Rate Readjustments
The objective of our optimization problem is to minimize power usage while maintaining the target rate requirements for all users in each network slice. When transmission power requirements exceed the BS available power due to high target rates, it is necessary to readjust the target rates of users in the CL slice to ensure that high-priority tasks in the URLLC and TS slices continue without interruption.
Each user is associated with a Lagrange multiplier, $\lambda_{i}$ during optimization as seen algorithm 1. $\boldsymbol{\lambda}$ indicates how much the overall power consumption will change if the rate constraints are relaxed or tightened. We employ a rate readjustment strategy [16] detailed in algorithm 2, to proportionally lower the target rates across users in the CL slice using their multipliers. The algorithm adopts the well known Newton-Raphson method [17] to iteratively calculate the difference between available and required power based on $\boldsymbol{\lambda}$, efficiently reducing rates step by step. The adjustments continue until power consumption is just below the maximum allowable limit, managed by a tolerance value .
The proportional rate reduction $\Delta r_{o,i}$ for each CL slice user $i$ is given by
$\Delta r_{o,i}=(p_{\text{opt}}-1)\left(\frac{\lambda_{i}}{\sum_{i=1}^{N}\lambda_{i}}\right),\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{C}$ (18)
where $p_{\text{opt}}=p_{\text{required}}/p_{\text{available}}$. The rate update formula is thus given by
$r_{o,i}=\max(r_{o,i}-\Delta r_{o,i},0),\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{C}$ (19)
By updating the rates with (19), users in CL slice contributing more to the power excess (higher $\lambda_{i}$) will experience a larger target rate reduction. Thereby efficiently reducing the total power by targeting the reductions more strategically based on users contributions to the power overshoot.
Algorithm 2 Rate Readjustment Algorithm
1:Input: $r_{o,i}$ $\forall i\in\mathcal{N}_{s},\forall s\in\mathcal{S}_{C}$, and $p_{\text{available}}$
2:Output: $r_{o,i}$ $\forall i\in\mathcal{N}_{s},\forall s\in\mathcal{S}_{C}$, and $p_{ij}$
3:Solve optimization problem (7) and obtain:
4:$p_{ij}$, and $\lambda_{i}$ $\forall i\in\mathcal{N}_{s}$, $\forall s\in\mathcal{S}_{C}$
5:Evaluate $p_{\text{required}}=\sum p_{ij}$, and $p_{\text{opt}}=\frac{p_{\text{required}}}{p_{\text{available}}}$
6:while $p_{\text{opt}}>1+\epsilon$ do
7:     $\Delta r_{o,i}=(p_{\text{opt}}-1)\left(\frac{\lambda_{i}}{\sum_{i=1}^{N}\lambda_{i}}\right)$
8:     $r_{o,i}=\max(r_{o,i}-\Delta r_{o,i},0)$
9:     Solve optimization problem (7) and obtain:
10:      $p_{ij}$, and $\lambda_{i}$ $\forall i\in\mathcal{N}_{s}$, $\forall s\in\mathcal{S}_{C}$
11:     Evaluate $p_{\text{required}}=\sum p_{ij}$, and $p_{\text{opt}}=\frac{p_{\text{required}}}{p_{\text{available}}}$
12:end while
@@ IV-A Scenario Description
For our simulation, we consider an indoor smart factory environment, characterised by densely packed reflecting elements of diverse shapes and sizes, and different user terminals each with varying QoS requirements that generate different traffic patterns. We consider a 5G BS deployed in this environment with a coverage area radius of 100 m having three slices, one each of CL, URLLC and TS slices. We use the 3GPP [18] non-line-of-sight (NLOS) path loss model for indoor factory (InF) dense clutter, low BS (InF-DL) scenario given in (20) in combination with Rayleigh fading, to obtain the channel gain in the links between the users and the BS
@@
Slice Requirements Target Rates No. of Active Users at Time Slot t
$t<33$ $33\leq t\leq 66$ $t>66$
CL $C_{s}=27\,\text{Mbps}$ 27 Mbps 5 2 7
URLLC $J_{s}=1\,\text{ms}\newline \gamma_{s}=99.9\%\newline a_{s}=2\,\text{Mbps}\newline D_{s}^{\text{max}}=10\,\text{ms}$ 2 Mbps 2 2 2
TS $L=5\,\text{kB}\newline T_{sp}=10\,\text{ms}$ 1.64 Mbps 1 1 1
@@ IV-B Slice Isolation and QoS Assessments
To evaluate the achievement of effective slice isolation, we show the mean user rates across the three slices in Fig.2(a). From this figure, we can observe that the mean rate of users in the CL slice experiences abrupt increases and decreases in response to the sudden changes in the number of users, decreasing and increasing respectively, at those specific times. However, the mean rate for the users in the URLLC and TS slices remains consistent, indicating complete isolation between the slices.
Given that all QoS constraints where translated into rate constraints, meeting the target rates for each slice equates to fulfilling their respective QoS requirements. Fig.2(b) shows the sum rates of users in each slice. This figure clearly indicates that all the target rates per slice, as outlined in Table II, have been met, thereby demonstrating the achievement of the delay, jitter, reliability and bitrate constraints for all users. An initial transient for the sum rate in the CL slice is observed as a result of the transient behaviour in the control loop in (17).
(a) Mean rate of users per slice
(b) Sum rate of users per slice
Figure 2: Achievement of (a) slice isolation and (b) target QoS for the three slices considered
@@ IV-C Admission Control
Slice New Target Rates No. of Active Users at Time Slot t
$t<33$ $33\leq t\leq 66$ $t>66$
CL 270 Mbps 10 7 12
URLLC 20 Mbps 4 4 4
TS 1.64 Mbps 2 2 2
Fig.3(a) shows the total BS power without admission control. We can observe that the power required to satisfy the QoS for all users across the three slices exceeds the available 23 dBm even when the number of users in the CL slice decreased at $33\leq t\leq 66$. This insufficiency indicates that the requirements of all active users will not be met, which can disrupt critical transmissions. To prevent this, we implement admission control to limit the required power to 23 dBm by selectively reducing the rates of users in the CL slice with less stringent requirements. The effects of this strategy are depicted in Fig.3(b) and Fig.4.
The total BS power when admission control is active is shown in Fig.3(b). The power requirement remains within the available limit of 23 dBm, confirming the success of our strategy. The quick drop in the total power observed at $t=33$ corresponds to the sudden decrease in the number of users at that time slot. Additionally, Fig.4(a) illustrates the rate adjustments in the CL slice during admission control. Specifically, at $t<33$ when the required power initially exceeds 23 dBm, the sum rate of the CL slice is reduced from the target of 270 Mbps to approximately 259 Mbps. At $33\leq t\leq 66$, as the number of users in the CL slice decreases from 10 to 7, the sum rate is readjusted to around 267 Mbps. Furthermore, when the number of users increased from 7 to 12 at $t>66$, the sum rate reduced again to 251 Mbps.
@@
Figure 3: BS transmit power (a) without and (b) with admission control
Figure 4: Sum rates (a) and mean rates (b) of users per slice during admission control
@@ V Conclusion and Future work
This research lays the groundwork for developing a dynamic, efficient, and autonomous slice-aware radio resource management system for smart factories. Future work will delve into the slicing resource allocation challenges posed by mobile terminals like AGV and evaluate performance using realistic industrial scenarios with actual factory channel traces and its application to uplink communications.
@@ VI Appendix I
Take a sequence $\mu_{1}\leq\mu_{2}\leq\dots\leq\mu_{N}$ of real values, and define $F(\mu)$ a function to be minimized with respect to $\mu$